### v2

# Perturbation Theory and the Quantum Rabi-model

Abstract.
In the first part of the paper we study a perturbative model of the Rabi system of Quantum Optics. We are therefore able to describe, through Rellich’s theory, an analytic expansion of finite families of eigenvalues, of arbitrary fixed length. In particular, we prove that for finite families of eigenvalues the Braak conjecture holds. In the second part we study the asymptotics of the Weyl spectral counting function of a class of systems that generalize the Quantum Rabi Model to an $N$-level atom ($N\geq 3$) with N1N-1 cavity modes of the electromagnetic field.
The authors are members of the Research Group GNAMPA of INdAM
The first author acknowledges the support by EPSRC Early Career Fellowship: EP/V001760/1.
The second author was partially supported by the Italian Ministry of University and Research, under PRIN2022 (Scorrimento) “Anomalies in partial differential equations and applications”, 2022HCLAZ8_002, J53C24002560006.
2000 Mathematics Subject Classification. Primary 35P20; Secondary 35S05, 81Q10
Key words and phrases: Spectral theory and eigenvalue problems for PDEs; Non-commutative harmonic oscillators; Weyl-calculus; Rabi system; Braak conjecture.
1. Introduction
The Quantum Rabi model (QRM), one of the most important models in Quantum Optics, describes the interaction of a 2-level atom and one cavity-mode electromagnetic field even when the field is not near resonance with the atomic transition and the coupling strength is not weak. In fact, it can be seen as the model leading to the Jaynes-Cummings model by rotating waves approximation, which is valid if the field is near resonance with the atomic transition, and the coupling strength is weak. (See Braak [2] and Rabi [20, 21].)
QRM is written as (with $\boldsymbol{a}$ the annihilation operator, $\boldsymbol{a}^{\dagger}$ the creation operator and with $I_{2}$ denoting the $2\times 2$ identity matrix)
(1.1) $Q_{\mathrm{Rabi}}=\boldsymbol{a}^{\dagger}\boldsymbol{a}I_{2}+\alpha(\boldsymbol{a}^{\dagger}+\boldsymbol{a})\left[\begin{array}[]{cc}0&1\\ 1&0\end{array}\right]+\Delta\left[\begin{array}[]{cc}1&0\\ 0&-1\end{array}\right],$
where $2\Delta$ is the splitting level (i.e. the difference in the energy of the atomic levels) and $\alpha>0$ describes the strength of the interaction between the photon and the atom.
Being a system on $\mathbb{R}^{n}$ with polynomial coefficients, based on the Quantum Harmonic Oscillator, it fits in the framework of Non-Commutative Harmonic Oscillators (NCHO) introduced by Wakayama and Parmeggiani in [18, 19] and as later extended in Parmeggiani [16, 17], and more recently in Malagutti and Parmeggiani [13] (and [14]). However, a difficulty in the geometric analysis of the Rabi system is that it is not a semiregular globally elliptic system in the sense of [13], in that, although the symbol is principally scalar, the symbol of the first order part (the semiprincipal symbol) cannot be smoothly diagonalized in the phase space $\mathbb{R}^{n}\times\mathbb{R}^{n}$ (as instead is the case for the Jaynes-Cummings system).
The paper is divided into two parts, both of which are perturbative in nature. In the first part we study a perturbation of the classical $1$-d QRM, which can be dealt with by the Rellich perturbation theory as in [22] (we chose that approach in place of Kato’s one [9] because more direct). In this part we destroy the homogeneity of the system, which allows us to bring the system into an isometric system to which we may apply perturbation theory. Since all the eigenvalues of the unperturbed system (a scalar quantum harmonic oscillator acting on $L^{2}(\mathbb{R};\mathbb{C}^{2})$) are degenerate, we apply Rellich’s approach by assuming some first-order (in the perturbation parameter) nondegeneracy, described by the way the interaction parameter is approximated by zeros of Legendre polynomials. That allows to control finite segments of the spectrum of the system, for a sufficiently small perturbation parameter, and to prove that in this context Braak’s Conjecture for $Q_{\mathrm{Rabi},\varepsilon}$, a particular case of our system $Q_{R}$ (see Section 5) holds, at least for arbitrarily long but finite parts of the spectrum. In addition, Rellich’s method gives control on the corresponding eigenfunctions. We also deal with ”degenerate” (in the sense of Rellich) situations, but in that case we use the simpler method of quasimodes (see Lazutkin [11]) that, however, gives up control on the eigenfunctions but retains control on the spectrum, proving its simplicity everywhere. Since in general we do not control in this case the eigenfunctions, we do not have information on the parity, which is basic for Braak’s conjecture. In all cases, our method, depending on the nondegeneracy of $Q^{(2)}_{N}$ (see Section 6), gives that by switching on the perturbation parameter $\varepsilon$ the perturbed eigenvalues may move to the right, or to the left, or both sides of the unperturbed eigenvalue $N$ (at $\varepsilon=0$) according to the signs of the eigenvalues of $Q^{(2)}_{N}$. This gives instances in which Braak’s conjecture keeps holding, however leaving open the possibility that in these degenerate cases Braak’s might no longer be true for very particular values of the interaction parameter. It also indicates a strategy to cope with more degenerate situations. In a sense, our result in the nondegenerate case may be thought of as a complement of Rudnik’s result [23]. For generalizations of the Braak’s Conjecture to asymmetric quantum Rabi models, see the recent paper by Braak, Nguyen, Reyes-Bustos and Wakayama [3].
In the second part of the paper we describe in the first place some possible generalizations of the QRM to higher dimensions for an interaction with atoms having several energy levels, where the space dimension $n$ is related to the number $N$ of atomic levels by the equation n=N1.n=N-1. In order to approach the study of the asymptotics of the Weyl spectral counting function, we have to cope with the problem of the non-diagonalizability in the symbol class of the first order part of the system. We then use a perturbation method that brings the perturbed symbol into the SMGES class, for which we may exploit a pseudodifferential diagonalization of the system and the results of [13]. Variational considerations then allow us to give a two-term asymptotics of the Weyl function. The result is not known in the higher dimensional and energy level number case.
We end this introduction by giving the plan of the paper.
In Section 2 we describe the Quantum Rabi Model from a physical perspective, highlighting its connection with the Jaynes-Cummings model by Rotating Wave Approximation and its importance as a foundational model in Quantum Optics. In Section 3 we recall our setup, and review the global pseudodifferential calculus used especially in the final Section 7. In Section 4 we give the perturbative result, under a non-degeneration condition, on finite initial segments of the spectrum of the QRM. In Section 5 we recall Braak’s conjecture and show its validity in our setting. We next deal with the degenerate case by the use of quasi-modes, which is the content of Section 6. We will show instances in which Braak’s conjecture keeps holding also in the degenerate case. In the final Section 7 we introduce the generalizations of the QRM and give a perturbative proof of a two-term asymptotics of the Weyl spectral counting function.
2. Physics of the Quantum Rabi Model
The Quantum Rabi Model (QRM) describes the fundamental interaction between a two-level quantum system (a qubit, atom, or spin-½) and a single bosonic mode (a quantized harmonic oscillator, typically a cavity or vibrational mode). It plays a central role in Quantum Optics, cavity and circuit QED, trapped-ion system, and modern quantum technologies, as it captures the minimal setting in which light-matter interaction is in full treated quantum mechanically.
2.1. Hamiltonian and basic structure
A general form for the Hamiltonian of the Quantum Rabi Model is
(2.2) $H=\hbar\omega\,\boldsymbol{a}^{\dagger}\boldsymbol{a}+\frac{\hbar\Omega}{2}\sigma_{z}+\hbar g\,\sigma_{x}(\boldsymbol{a}+\boldsymbol{a}^{\dagger}),$
where $\omega$ is the frequency of the bosonic mode, $\Omega$ is the energy splitting of the two-level system, $g$ is the coupling strength, $\boldsymbol{a},\boldsymbol{a}^{\dagger}$ are the bosonic annihilation and creation operators, and $\sigma_{x},\sigma_{z}$ are Pauli matrices acting on the two-level system, that is
$\sigma_{x}=\left[\begin{array}[]{cc}0&1\\ 1&0\end{array}\right],\quad\sigma_{y}=\left[\begin{array}[]{cc}0&-i\\ i&0\end{array}\right],\quad\sigma_{z}=\left[\begin{array}[]{cc}1&0\\ 0&-1\end{array}\right],$
Writing $\sigma_{\pm}=(\sigma_{x}\pm i\sigma_{y})/2$, the interaction term becomes
(2.3) $H_{\text{int}}=\hbar g(\sigma_{+}\boldsymbol{a}+\sigma_{-}\boldsymbol{a}^{\dagger}+\sigma_{+}\boldsymbol{a}^{\dagger}+\sigma_{-}\boldsymbol{a}),$
where the first two terms correspond to excitation exchange (rotating terms) and the latter two correspond to simultaneous creation or annihilation of excitations (counter-rotating terms). Neglecting the counter-rotating terms yields the Jaynes–Cummings model (as obtained by Jaynes and Cummings in [8], see also the book by Scully and Zubairy [24]).
2.2. Symmetry and integrability
Although the QRM does not preserve excitation number, it possesses a discrete $\mathbb{Z}_{2}$ parity symmetry given by
(2.4) $\Pi=\sigma_{z}(-1)^{\boldsymbol{a}^{\dagger}\boldsymbol{a}},$
which commutes with the Hamiltonian. This symmetry splits the Hilbert space into even and odd parity subspaces and is essential for understanding the structure of the spectrum. Braak [1] showed that the QRM is integrable in a generalized sense and that its spectrum can be obtained from the zeros of transcendental functions having no simple closed-form expression, but built from infinite series and special functions.
2.3. Coupling regimes
The physics of the QRM depends strongly on the ratio of the coupling strength to the bare frequencies.
2.3.1. Weak coupling: Jaynes-Cummings regime
For $g\ll\omega,\Omega$, and near resonance $\omega\approx\Omega$, the rapidly oscillating counter-rotating terms can be neglected under the rotating-wave approximation (RWA). The Hamiltonian reduces to
(2.5) $H_{\text{JC}}=\hbar\omega\boldsymbol{a}^{\dagger}\boldsymbol{a}+\frac{\hbar\Omega}{2}\sigma_{z}+\hbar g(\sigma_{+}\boldsymbol{a}+\sigma_{-}\boldsymbol{a}^{\dagger}),$
which preserves the total excitation number, as showed by Jaynes and Cummings in [8].
2.3.2. Ultrastrong coupling
When $g/\omega=0.1\pm\epsilon$, the counter-rotating terms become relevant and the RWA breaks down. This is the ultrastrong coupling regime, experimentally realized in superconducting circuit QED (see the paper by Niemczyk [15]). The ground state contains virtual photons and the system exhibits phenomena such as the Bloch–Siegert shift.
2.3.3. Deep strong coupling
For $g/\omega\gtrsim 1$, the system enters the deep strong coupling regime [5]. In this limit the interaction term dominates and the eigenstates are highly entangled superpositions of qubit and oscillator states.
2.4. Physical interpretation
The QRM describes the dressing of the two-level system by the quantized field. In weak coupling this dressing leads to small energy shifts, while in strong coupling it produces hybrid light-matter excitations (polaritons). The counter-rotating terms allow virtual excitations, which influence measurable quantities such as spectra and ground-state photon populations (for a more complete description of the topic see Niemczyk, Deppe, Huebl et al. [15], and the survey by Forn–Díaz, García-Ripoll, Peropadre et al. [5]).
2.5. Experimental realizations
Flexible realizations of the QRM exist in superconducting circuit QED, trapped ions, and optomechanical systems, where parameters can be tuned to explore ultrastrong and deep strong coupling regimes [15, 5].
3. The setup
In this section we describe our setting. We deal with globally elliptic systems (see e.g. Helffer [6] or Shubin [25], or Parmeggiani [16]) with polynomial coefficients.
We work in the phase-space $T^{\ast}\mathbb{R}^{n}=\mathbb{R}^{n}\times\mathbb{R}^{n}=\mathbb{R}^{2n}$, writing $X=(x,\xi)\in\mathbb{R}^{2n}$ for the points in $\mathbb{R}^{2n}$. We write for short $\dot{\mathbb{R}}^{2n}=\mathbb{R}^{2n}\setminus\{0\}$. As usual, we put $\langle X\rangle=(1+|X|^{2})^{1/2}$, $D_{x_{j}}=-i\partial_{x_{j}}$. The $L^{2}$-inner product in $L^{2}(\mathbb{R}^{n})$ (or $L^{2}(\mathbb{R}^{n};\mathbb{C}^{N})$) is denoted by $(\cdot,\cdot)_{0}$.
We next recall the Shubin calculus of global operators (in the context of the Weyl-Hörmander calculus).
Definition 3.1.
Let $a\in C^{\infty}(\mathbb{R}^{2n})$ and $m\in\mathbb{R}$. We say that $a$ is a global symbol of order $m$, and write $a\in S^{m}(\mathbb{R}^{n})$, if for all $\alpha\in\mathbb{Z}_{+}^{2n}$ there exists $C_{\alpha}>0$ such that
(3.1) $|\partial_{X}^{\alpha}a(X)|\leq C_{\alpha}\langle X\rangle^{m-|\alpha|},\quad X\in\mathbb{R}^{2n}.$
A symbol $a\in S^{m}(\mathbb{R}^{n})$ ($m\geq 0$) is elliptic when there is $c_{0}>0$ such that
$|a(X)|\geq c_{0}\langle X\rangle^{m},\quad\forall X.$
In the Hörmander-Weyl calculus (see Hörmander [7]), $S(\langle\cdot\rangle^{m};g)=S^{m}(\mathbb{R}^{n})$ where the Hörmander metric is given by
$X\longmapsto g_{X}=\frac{|dX|^{2}}{\langle X\rangle^{2}}.$
We denote
$S^{\infty}(\mathbb{R}^{n})=\bigcup_{m\in\mathbb{R}}S^{m}(\mathbb{R}^{n})$
and
$S^{-\infty}(\mathbb{R}^{n})=\bigcap_{m\in\mathbb{R}}S^{m}(\mathbb{R}^{n})=\mathscr{S}(\mathbb{R}^{2n}).$
If $a\in S^{m}(\mathbb{R}^{n})$, we associate with it a pseudodifferential operator using the so-called Weyl-quantization.
Definition 3.2.
Let $a\in S^{m}(\mathbb{R}^{n})$. The Weyl-quantized pseudodifferential operator associated with $a$ is defined as
$a^{\mathrm{w}}(x,D)u(x)=(2\pi)^{-n}\iint_{\mathbb{R}^{2n}}e^{i\langle x-y,\xi\rangle}a((x+y)/2,\xi)u(y)dyd\xi,\quad u\in\mathscr{S}(\mathbb{R}^{n}).$
One has that $a^{\mathrm{w}}(x,D)$ is a linear operator that acts continuously $\mathscr{S}(\mathbb{R}^{n})\to\mathscr{S}(\mathbb{R}^{n})$ and extends by duality to a linear continuous operator $\mathscr{S}^{\prime}(\mathbb{R}^{n})\to\mathscr{S}^{\prime}(\mathbb{R}^{n})$.
Definition 3.3.
We define the set of global pseudodifferential operators of order $m$ as
$\Psi^{m}(\mathbb{R}^{n}):=\Bigl\{A\colon\mathscr{S}^{\prime}(\mathbb{R}^{n})\xrightarrow[\textit{\tiny cont.}]{\textit{\tiny linear}}\mathscr{S}^{\prime}(\mathbb{R}^{n});\ \exists\,a\in S^{m}(\mathbb{R}^{n}),\;A=a^{\mathrm{w}}(x,D)\Bigr\}.$
We write
$\Psi^{\infty}(\mathbb{R}^{n})=\bigcup_{m\in\mathbb{R}}\Psi^{m}(\mathbb{R}^{n}),\quad\Psi^{-\infty}(\mathbb{R}^{n})=\bigcap_{m\in\mathbb{R}}\Psi^{m}(\mathbb{R}^{n}).$
We call the elements of $\Psi^{-\infty}(\mathbb{R}^{n})$ smoothing operators: they are those operators in $\Psi^{\infty}(\mathbb{R}^{n})$ that have smoothing, i.e. $\mathscr{S}(\mathbb{R}^{n}\times\mathbb{R}^{n})$, Schwartz kernel.
The formal adjoint of $a^{\mathrm{w}}(x,D)$ is simply ${\bar{a}}^{\mathrm{w}}(x,D)$ and the composition law is written as follows: If $a\in S^{m}(\mathbb{R}^{n}$ and $b\in S^{m^{\prime}}(\mathbb{R}^{n})$ then $a^{\mathrm{w}}(x,D)b^{\mathrm{w}}(x,D)=(a\sharp b)^{\mathrm{w}}(x,D)$ where $a\sharp b\in S^{m+m^{\prime}}(\mathbb{R}^{n})$ is written, for any given $N\in\mathbb{Z}_{+}$, as
$a\sharp b(X)=\sum_{k=0}^{N}\frac{1}{k!}\left(\frac{i}{2}\sigma(D_{X};D_{Y})\right)^{k}a(X)b(Y)\bigl|_{X=Y}+r_{N+1}(X),$
with $r_{N+1}\in S^{m+m^{\prime}-(N+1)}(\mathbb{R}^{n})$ and where
$\sigma(D_{X};D_{Y})=\sigma(D_{x},D_{\xi};D_{y},D_{\eta})=\langle D_{\xi},D_{y}\rangle-\langle D_{\eta},D_{x}\rangle.$
We recall also the important class $S^{m}_{\mathrm{sreg}}(\mathbb{R}^{n})$ of semiregular symbols.
Definition 3.4.
We say that $a\in S^{m}(\mathbb{R}^{n})$ belongs to the class $S^{m}_{\mathrm{sreg}}(\mathbb{R}^{n})$ of semiregular symbols, if $a$ possesses an asymptotic expansion
$a\sim\sum_{j\geq 0}a_{m-j},$
where each $a_{m-j}$ is homogeneous in $X\in\mathbb{R}^{n}\times\mathbb{R}^{n}\setminus\{(0,0)\}$ of degree mj,m-j, that is
$a_{m-j}(tX)=t^{m-j}a_{m-j}(X),\quad\forall t>0,\,\,\,\forall X\not=0.$
Asymptotic means that, chosen an excision function $\chi$ (i.e. $\chi\in C^{\infty}(\mathbb{R}^{2n}),$ $0\leq\chi\leq 1,$ $\chi\equiv 0$ when, say, $|X|\leq 1/2,$ $\chi\equiv 1$ when, say, $|X|\geq 1$), for every $N\in\mathbb{Z}_{+}$ we have
$a-\chi\sum_{j=0}^{N}a_{m-j}\in S^{m-(N+1)}(\mathbb{R}^{n}).$
The terms $a_{m},$ $a_{m-1},$ $a_{m-2}$ and $a_{m-3}$ are, respectively, called the principal, semiprincipal, subprincipal, semisubprincipal symbols of the operator $a^{\mathrm{w}}(x,D).$
A fundamental feature of the Hörmander-Weyl calculus is its invariance under the action of the affine symplectic group (see Hörmander [7]). Given $\chi\colon\mathbb{R}^{n}\times\mathbb{R}^{n}\longrightarrow\mathbb{R}^{n}\times\mathbb{R}^{n}$ an affine symplectomorphism, there exists a metaplectic operator $U_{\chi}$, unitary in $L^{2}$ and automorphisms of both $\mathscr{S}$ and $\mathscr{S}^{\prime}$ (uniquely determined modulo a unit complex number) such that
(3.2) $U_{\chi}^{*}a^{\mathrm{w}}(x,D)U_{\chi}=(a\circ\chi)^{\mathrm{w}}(x,D).$
It is important to recall that semiregular symbols are preserved by the action of the linear symplectic group (as follows from (3.2)).
[…]
$S^{m}(\mathbb{R}^{n};\mathsf{M}_{N})=S^{m}(\mathbb{R}^{n})\otimes\mathsf{M}_{N},$
where $\mathsf{M}_{N}$ denotes the space of the $N\times N$ complex matrices. The composition formulas are then to be understood as compositions rows-times-columns, and the adjoint of a symbol will be given by the adjoint matrix of symbolic entries. Ellipticity is defined by requiring that $\det a\in S^{mN}(\mathbb{R}^{n})$ is elliptic. Analogously, one has the class of matrix-valued semiregular symbols $S^{m}_{\mathrm{sreg}}(\mathbb{R}^{n};\mathsf{M}_{N})$. In this case, ellipticity is a condition determined only by the principal symbol.
It is convenient to recall the definition of the class of SMGES semiregular systems as introduced in [13] (Def. 2.4).
Definition 3.5.
A semiregular symbol $a=a^{*}\in S^{m}_{\mathrm{sreg}}(\mathbb{R}^{n};\mathsf{M}_{N})$ is said to be a semiregular metric globally elliptic system (SMGES) if it is principally scalar with a globally elliptic principal part, and whose semiprincipal part is smoothly diagonalizable by unitary elements of $S^{0}_{\mathrm{sreg}}(\mathbb{R}^{n};\mathsf{M}_{N})$ (for $X\not=0$) into distinct blocks that can be ordered in strictly ascending order for all $X\not=0$.
Next, as regards the $1$-d QRM we have the following setup. Given $p_{2}(X)=|X|^{2}/2$, $X\in\mathbb{R}^{2}$, we have that $P_{0}=p_{2}^{\mathrm{w}}(x,D)\in\Psi^{2}(\mathbb{R})$ is the usual quantum harmonic oscillator $(D_{x}^{2}+x^{2})/2,$ $x\in\mathbb{R}.$ We have that $P_{0}$, realized as an unbounded operator in $L^{2}(\mathbb{R})$ with maximal domain the Shubin Sobolev space $B^{2}(\mathbb{R})$ (see [6, 25]), is self-adjoint and has a discrete spectrum made of eigenvalues with multiplicity $1$ given by the sequence $\{N+1/2\}_{N\in\mathbb{Z}_{+}}$, and corresponding eigenfunctions, the ones we will use in our subsequent computations, given by the Schwartz functions
(3.3) $\varphi_{N}(x)=\left(\frac{x-\partial_{x}}{\sqrt{2}}\right)^{N}e^{-x^{2}/2}=H_{N}(x)e^{-x^{2}/2},\quad N\in\mathbb{Z}_{+}$
(the $H_{N}$ being the Hermite polynomials of degree $N$), that form an orthogonal basis of $L^{2}(\mathbb{R})$ (and of $\mathscr{S}(\mathbb{R})$ and of $\mathscr{S}^{\prime}(\mathbb{R})$). One has
$(\varphi_{N},\varphi_{N^{\prime}})_{0}=\sqrt{\pi}\,N!\delta_{NN^{\prime}},\quad N,N^{\prime}\in\mathbb{Z}_{+}.$
Recall also, in passing, an important property of the Hermite polynomials that will be needed shortly:
(3.4) $H_{N}(x+y)=\sum_{k=0}^{N}\binom{N}{k}H_{k}(x)(2y)^{N-k},\quad\forall x,y\in\mathbb{R}.$
The Rabi system we consider in this paper is the following. Let $\alpha,\gamma_{1},\gamma_{2}\in\mathbb{R}$ with $\alpha\not=0$ and $\gamma_{1}>\gamma_{2}$. For $\varepsilon\in(-1,1)$ we consider the system ($I_{2}$ being the $2\times 2$ identity)
(3.5) $Q_{R}=Q_{R,\varepsilon}(x,D)=p_{2}^{\mathrm{w}}(x,D)I_{2}+\alpha x\left[\begin{array}[]{cc}0&1\\ 1&0\end{array}\right]+\varepsilon\left[\begin{array}[]{cc}\gamma_{1}&0\\ 0&\gamma_{2}\end{array}\right],$
with maximal domain
$D_{R}=\{u\in L^{2}(\mathbb{R};\mathbb{C}^{2});\,\,Q_{R}u\in L^{2}(\mathbb{R};\mathbb{C}^{2})\}=B^{2}(\mathbb{R};\mathbb{C}^{2})$
(hence, it does not depend on $\varepsilon$). The operator $Q_{R}$, thought of as an unbounded operator on $L^{2}$ with domain $D_{R}$ is then self-adjoint with a compact resolvent. In fact, its symbol $Q_{R}(X)$ belongs to $\Psi^{2}(\mathbb{R};\mathsf{M}_{2})$ and is globally elliptic, i.e. the principal part of degree $2$ is invertible for $X\not=0$. By the compact embedding of $D_{R}$ into $L^{2}$, the resolvent operator is therefore compact so that $Q_{R}$ has a real and discrete spectrum made of an increasing sequence of eigenvalues with finite multiplicities. As we will see, by Rellich’s theory, one has that its spectrum can be parametrized by analytic functions of $\varepsilon$ (see Kato [9], Thm. 3.9 page 392–393).
Define next the $L^{2}$-isometries
$T_{\pm}\colon L^{2}(\mathbb{R})\longrightarrow L^{2}(\mathbb{R}),\quad u\longmapsto u(\cdot\mp\alpha),$
that are metaplectic operators associated with the symplectomorphisms $\chi_{\pm}\colon(x,\xi)\longmapsto(x\pm\alpha,\xi)$. Then, since $T_{\pm}=T_{\mp}^{*}$, we have
$T_{\mp}p_{2}^{\mathrm{w}}(x,D)T_{\pm}=(D^{2}+(x\pm\alpha)^{2})/2.$
The following lemma is fundamental.
Lemma 3.6.
There exists a metaplectic operator $U\colon L^{2}(\mathbb{R};\mathbb{C}^{2})\longrightarrow L^{2}(\mathbb{R};\mathbb{C}^{2})$ (which is an automorphism of $\mathscr{S}(\mathbb{R};\mathbb{C}^{2})$ and of $\mathscr{S}^{\prime}(\mathbb{R};\mathbb{C}^{2})$) such that
(3.6) $U^{*}(Q_{R,\varepsilon}+\frac{\alpha^{2}}{2}I_{2})U=A_{\varepsilon}=A+\varepsilon B=p_{2}^{\mathrm{w}}(x,D)I_{2}+\varepsilon\left[\begin{array}[]{cc}\beta_{1}&\beta_{2}T_{+}^{2}\\ \beta_{2}T_{-}^{2}&\beta_{1}\end{array}\right],$
where
$\beta_{1}=(\gamma_{1}+\gamma_{2})/2,\quad\beta_{2}=(\gamma_{1}-\gamma_{2})/2.$
Proof.
Let $U_{0}=\left[\begin{array}[]{cc}1&1\\ 1&-1\end{array}\right]/\sqrt{2},$ and $T=\left[\begin{array}[]{cc}T_{+}&0\\ 0&T_{-}\end{array}\right].$ Then
$U_{0}^{*}(Q_{R,\varepsilon}+\frac{\alpha^{2}}{2}I_{2})U_{0}=\left[\begin{array}[]{cc}\frac{1}{2}(D^{2}+(x+\alpha)^{2})&0\\ 0&\frac{1}{2}(D^{2}+(x-\alpha)^{2})\end{array}\right]+\varepsilon\left[\begin{array}[]{cc}\beta_{1}&\beta_{2}\\ \beta_{2}&\beta_{1}\end{array}\right]$
$T^{*}\Bigl(p_{2}^{\mathrm{w}}(x,D)I_{2}+\varepsilon T\left[\begin{array}[]{cc}\beta_{1}&\beta_{2}\\ \beta_{2}&\beta_{1}\end{array}\right]T^{*}\Bigr)T,$
which proves the lemma with $U=U_{0}T^{*}$. ∎
Remark 3.7.
Note that the symbol $t_{\pm}$ of $T_{\pm}$ is given by $t_{\pm}(X)=e^{\pm ix\xi}$ which belongs to $S(1,|dX|^{2}).$ In particular, for the composition $S(1;|dX|^{2})\sharp S^{m}(\mathbb{R})$ we just have
$S(1;|dX|^{2})\sharp S^{m}(\mathbb{R})\subset S(\langle\cdot\rangle^{m};|dX|^{2}).$
Hence there is no asymptotic calculus, but the continuity properties (on $L^{2}$ and on Shubin’s Sobolev spaces, see [6]) keep holding true.
3.1. Relation between $Q_{\mathrm{Rabi},\varepsilon}$ and $Q_{R}$
Strictly speaking, the Rabi system we consider here is a translation by $\frac{1}{2}I_{2}$ of the ”genuine” Rabi system (1.1), that we may write (replacing $\Delta$ in (1.1) by $\varepsilon\Delta$, and putting $g=\alpha/\sqrt{2}$) as
$Q_{\mathrm{Rabi},\varepsilon}=p_{2}^{\mathrm{w}}(x,D)I_{2}+\alpha x\left[\begin{array}[]{cc}0&1\\ 1&0\end{array}\right]+\varepsilon\Delta\left[\begin{array}[]{cc}1&0\\ 0&-1\end{array}\right]-\frac{1}{2}I_{2},$
whence
$Q_{R,\varepsilon}=Q_{\mathrm{Rabi},\varepsilon}+\frac{1}{2}I_{2},$
with $\gamma_{1}=-\gamma_{2}$ and $\gamma_{1}=\Delta$. Therefore up to translating the whole spectrum of $Q_{R}$ by $1/2$, studying the spectrum of $Q_{R,\varepsilon}$ with $\Delta=\gamma_{1}=-\gamma_{2}$ is the same as studying the spectrum of $Q_{\mathrm{Rabi},\varepsilon}$.
Using the metaplectic operator $U$ of Lemma 3.6 above, one has that
$U^{*}Q_{\mathrm{Rabi},\varepsilon}U=p_{2}^{\mathrm{w}}(x,D)I_{2}+\varepsilon\Delta\left[\begin{array}[]{cc}0&T_{+}^{2}\\ T_{-}^{2}&0\end{array}\right]-\frac{1+\alpha^{2}}{2}I_{2}.$
4. Perturbation Theory
[…]
because of the $L^{2}\to L^{2}$ continuity of $BE$ and $BR$. Therefore: The family $\varepsilon\mapsto A+\varepsilon B$ is a holomorphic regular family in the sense of Rellich [22] (equivalently, of type (A) in the sense of Kato) for all $|\varepsilon|\leq 1/(2C_{B})=:\varepsilon_{0},$ see [9]. We denote by $I_{\varepsilon_{0}}$ the corresponding interval centered at 0. It follows that, by the very nature of $A+\varepsilon B$, it has a compact resolvent for all $\varepsilon$ and hence its eigenvalues can be described as analytic functions of $\varepsilon$ (see Kato [9], Thm. 3.9 page 392–393). We may therefore study the analyticity of finite families of eigenvalues of $Q_{R,\varepsilon}$ in $\varepsilon$.
Since $A$ has spectrum $\{N+1/2;\,\,N\in\mathbb{Z}_{+}\}$ with constant multiplicity $2$, we have to prepare the ground for Rellich’s perturbation of degenerate eigenvalues. To that purpose, we have to study the quadratic form defined by $B$ restricted to the eigenspace of $A$ we are studying.
For some $N\in\mathbb{Z}_{+}$, let $\lambda_{0}=N+1/2$, $\varphi=\varphi_{N}$, and
[…]
where $e_{1},e_{2}$ is the canonical basis of $\mathbb{C}^{2}$. If $\pi_{E_{0}}\colon L^{2}(\mathbb{R};\mathbb{C}^{2})\longrightarrow E_{0}$ denotes the orthogonal projection onto $E_{0}$, we then have a bounded operator $R_{0}\colon L^{2}(\mathbb{R};\mathbb{C}^{2})\longrightarrow\mathrm{Ran}(A)\subset L^{2}(\mathbb{R};\mathbb{C}^{2})$ such that
$\left\{\begin{array}[]{ll}R_{0}(A-\lambda_{0})=(A-\lambda_{0})R_{0}=1-\pi_{E_{0}},\\ R_{0}\pi_{E_{0}}=0.\end{array}\right$.
(Therefore $R_{0}$ is a parametrix of the globally elliptic operator $A-\lambda_{0},$ because $\pi_{E_{0}}$ is a smoothing term since it has rapidly decreasing Schwartz kernel.)
Consider the bilinear form $(B\cdot,\cdot)_{0}\bigl|_{E_{0}\times E_{0}}$: for linear combinations $u=w_{1}\boldsymbol{\phi}_{1}+w_{2}\boldsymbol{\phi}_{2}\in E_{0}$, $w=\left[\begin{array}[]{c}w_{1}\\ w_{2}\end{array}\right]\in\mathbb{C}^{2}$, and likewise for $u^{\prime}\in E_{0}$ with coefficients $w^{\prime}=\left[\begin{array}[]{c}w^{\prime}_{1}\\ w^{\prime}_{2}\end{array}\right]$, we have
$(Bu,u^{\prime})_{0}=\langle Mw,w^{\prime}\rangle_{\mathbb{C}^{2}}$
($\langle\cdot,\cdot\rangle_{\mathbb{C}^{2}}$ being the Hermitian product in $\mathbb{C}^{2}$), where
[…]
(Note that $(T_{+}\varphi,T_{-}\varphi)_{0}\in\mathbb{R}$.) The eigenvalues of $M$ are given by
[…]
Then $\mu_{+}=\mu_{-}=\beta_{1}=(\gamma_{1}+\gamma_{2})/2$ (which is then independent of $N$) iff $\beta_{2}(T_{+}\varphi,T_{-}\varphi)_{0}=0$, that is, iff $(T_{+}\varphi,T_{-}\varphi)_{0}=0$. So, to have a nondegenerate splitting in the perturbed eigenvalue we have to study $(T_{+}\varphi,T_{-}\varphi)_{0}$ for a given Hermite eigenfunction $\varphi$ of $A$.
(We decided to keep the modulus of $(T_{+}\varphi,T_{-}\varphi)_{0}$ because then the perturbed eigenvalues $\lambda_{N,\pm}(\varepsilon)$ are ordered, for $\varepsilon$ sufficiently small depending on $N$.)
4.1. Study of $(T_{+}\varphi,T_{-}\varphi)_{0}$
We have the following lemma.
Lemma 4.1.
For all $N,k\in\mathbb{Z}_{+}$ the following formula holds
[…]
where
(4.2) $p_{N,k}(Z)=\sum_{j=1}^{\min\{N,k\}}(-1)^{k-j}\binom{N}{j}\binom{k}{j}j!Z^{N+k-2j}.$
In particular, for $N=k$,
(4.3) $(T_{+}\varphi_{N},T_{-}\varphi_{N})_{0}=e^{-\alpha^{2}}\sum_{j=0}^{N}(-4\alpha^{2})^{N-j}\binom{N}{j}^{2}|\!|\varphi_{j}|\!|_{0}^{2}=:\sqrt{\pi}\,e^{-\alpha^{2}}p_{N}(-4\alpha^{2}),$
where
(4.4) $p_{N}(Z)=\sum_{j=0}^{N}\binom{N}{j}^{2}(N-j)!Z^{j}=N!\sum_{j=0}^{N}\binom{N}{j}\frac{Z^{j}}{j!}.$
Proof.
In the first place notice that $\varphi_{N}|\!|_{0}^{2}=\sqrt{\pi}N!.$ Then
$(T_{+}\varphi_{N},T_{-}\varphi_{k})_{0}=e^{-\alpha^{2}}\int e^{-x^{2}}H_{N}(x-\alpha)H_{k}(x+\alpha)dx$
(by the Taylor expansion of Hermite polynomials (3.4))
$e^{-\alpha^{2}}\sum_{j=0}^{N}\sum_{j^{\prime}=0}^{k}\binom{N}{j}\binom{k}{j^{\prime}}(-2\alpha)^{N-j}(2\alpha)^{k-j^{\prime}}\int e^{-x^{2}}H_{j}(x)H_{j^{\prime}}(x)dx$
[…]
which concludes the proof. ∎
The next important observation is that
$p_{N}(-Z)=N!L_{N}(Z),$
where
$L_{N}(Z)=\sum_{j=0}^{N}(-1)^{j}\binom{N}{j}\frac{Z^{j}}{j!}$
is the $N$-th Laguerre polynomial. Therefore the vanishing of $(T_{+}\varphi,T_{-}\varphi)_{0}$ depends on the zeros of the Laguerre polynomials. It is known that the zeros of the $L_{k}$s are dense in $0,+\infty)$ (see Chihara [4]). Consider
$\mathsf{Z}:=\{x\in[0,+\infty);\,\,\exists k\in\mathbb{N},\,\,L_{k}(x)=0\}=\bigcup_{k\in\mathbb{N}}L^{-1}_{k}(0).$
Hence $\mathsf{Z}$ is countable and dense in $0,+\infty)$, whence $0,+\infty)\setminus Z$ is uncountable. We have that $L_{k}^{-1}(0)$ is given by
$0<x_{k,1}<x_{k,2}<\ldots<x_{k,k}.$
For a given $\delta>0$ let $I_{\delta}(x_{0})=\{x>0;\,\,|x-x_{0}|<\delta\}.$
Lemma 4.2.
Let $x_{0}\in[0,+\infty)\setminus\mathsf{Z}.$ There exists a sequence $\{k_{j}\}_{j\geq 1}\subset\mathbb{N}$ with $k_{j}\nearrow+\infty$ as $j\to+\infty$, and a sequence $\{\delta_{j}\}_{j\geq 1}\subset(0,+\infty)$ with $\delta_{j}\searrow 0$ as $j\to+\infty$, such that for every $j$, for all $k\leq k_{j}$ the polynomials $L_{k}$ have all the roots at distance at least $\delta_{j}$ from $x_{0}.$
Proof.
Let $\delta_{1}=1/10$ and
$k_{1}=\min\{k\in\mathbb{N};\,\,L_{k}^{-1}(0)\cap I_{\delta_{1}}(x_{0})\not$$\emptyset\}$.
Choose the least $\nu_{1}$, $1\leq\nu_{1}\leq k_{1}$, so that the root $x_{k_{1},\nu_{1}}$ of $L_{k_{1}}$ is closest to $x_{0}$ (of course, it cannot be $x_{0}$ by hypothesis). Put then
$\delta_{2}=\frac{|x_{k_{1},\nu_{1}}-x_{0}|}{10}<\frac{\delta_{1}}{10}=10^{-2},$
and define
$k_{2}=\min\{k>k_{1};\,\,L_{k}^{-1}(0)\cap I_{\delta_{2}}(x_{0})\not$$\emptyset\}$.
Choose then $\nu_{2},$ $1\leq\nu_{2}\leq k_{2}$, such that $x_{k_{2},\nu_{2}}$ is closest to $x_{0}.$ By induction we therefore have sequences $\{k_{j}\}_{j\geq 1},$ $\{\delta_{j}\}_{j\geq 1},$ and $\{\nu_{j}\}_{j\geq 1}$ such that $k_{j}\to+\infty,$ $x_{k_{j},\nu_{j}}\in L_{k_{j}}^{-1}(0)$ for all $j\geq 1,$
$\delta_{j+1}=\frac{|x_{k_{j},\nu_{j}}-x_{0}|}{10}<\frac{\delta_{j}}{10}<10^{-(j+1)}\searrow 0\,\,\text{\rm as}\,\,j\to+\infty,$
$x_{k_{j},\nu_{j}}\xrightarrow[j\to+\infty]{}x_{0},\,\,\text{\rm with}\,\,|x_{k_{j},\nu_{j}}-x_{0}|\searrow 0.$
This concludes the proof. ∎
In other words, Lemma 4.2 makes sure that if $x_{0}=4\alpha^{2}\in(0,+\infty)\setminus\mathsf{Z}$ then given any $j\geq 1$ we may find $k_{j}\geq 1$ such that for all integers $N\leq k_{j}$ the Laguerre polynomials $L_{N}$ do not have zeros at distance less than $\delta_{j}.$ Hence, by Lemma 4.1 for all such $N$ we have that $(T_{+}\varphi,T_{-}\varphi)_{0}\not=0$, whence (as we will see through Rellich’s theory) the perturbations of $\lambda_{N}(0)$ (which has multiplicity 2) split into simple eigenvalues $\lambda^{(N)}_{\pm}(\varepsilon)$. Recall, in fact, that
[…]
Using Satz 2, page 365 of Rellich [22] we therefore have proved the following result.
Theorem 4.3.
Let $\alpha$ be such that $4\alpha^{2}\in(0,+\infty)\setminus\mathsf{Z}$, and let $\{k_{j}\}_{j\geq 1}$ and $\{\delta_{j}\}_{j\geq 1}$ be the sequences of Lemma 4.2. Fix $k_{j}.$ Then, there exists $c_{j}\in(0,1)$ such that for all $N\leq k_{j}$ we have analytic functions
$(-c_{j},c_{j})\ni\varepsilon\mapsto\lambda^{(N)}_{\pm}(\varepsilon)\in\mathbb{R},$
$\lambda^{(N)}_{\pm}(\varepsilon)=\sum_{k=0}^{\infty}\lambda_{k,\pm}^{(N)}\varepsilon^{k},\quad\lambda_{0,\pm}^{(N)}=N+1/2,\,\,\,\lambda_{1,\pm}^{(N)}=\mu_{\pm},$
and
$(-c_{j},c_{j})\ni\varepsilon\mapsto u^{(N)}_{\pm}(\varepsilon)\in L^{2}(\mathbb{R};\mathbb{C}^{2})$
$u^{(N)}_{\pm}(\varepsilon)=\sum_{k=0}^{\infty}u_{k,\pm}^{(N)}\varepsilon^{k},\quad u^{(N)}_{0,\pm}=\boldsymbol{\phi}_{\pm}$
with $u^{(N)}_{\pm}(\varepsilon)\in D(A+\varepsilon B)$ for all $|\varepsilon|<c_{j}$ and
$u^{(N)}_{\pm}(\varepsilon)|\!|_{0}=1,\,\,\forall|\varepsilon|<c_{j},\quad(u^{(N)}_{0,\pm},u^{(N)}_{k,\pm})_{0}\in\mathbb{R},\,\,\forall k\geq 1,$
and with $u^{(N)}_{k,\pm}\in\mathscr{S}(\mathbb{R};\mathbb{C}^{2})$ for all $k$ and convergence of the series in $L^{2}$, such that
$(A+\varepsilon B)u^{(N)}_{\pm}(\varepsilon)=\lambda^{(N)}_{\pm}(\varepsilon)u^{(N)}_{\pm}(\varepsilon),\quad\forall|\varepsilon|<c_{j}.$
By unitary conjugation we therefore obtain also the following result for the eigenvalues of our Rabi quantum model $Q_{R}$.
Corollary 4.4.
In particular, for $N\leq k_{j}$ (and $|\varepsilon|<c_{j}$) the first $2(k_{j}+1)$ eigenvalues
$\mu_{\pm}^{(N)}(\varepsilon)=\lambda^{(N)}_{\pm}(\varepsilon)-\alpha^{2}/2,\quad 0\leq N\leq k_{j},$
of $Q_{R,\varepsilon}$ are all simple when $\varepsilon\not=0$.
We will call the $\lambda^{(N)}_{\pm}(\varepsilon)=\mu^{(N)}_{\pm}(\varepsilon)+\alpha^{2}/2$ the shifted eigenvalues of $Q_{R,\varepsilon}$.
Recall, at this point, the parity associated with Rabi’s system $Q_{\mathrm{Rabi},\varepsilon}$ (but also with our system $Q_{R,\varepsilon}$), given by the parity operator
$\Pi=(-1)^{\boldsymbol{a}^{\dagger}\boldsymbol{a}}\sigma_{z}.$
Since $\Pi$ commutes with $Q_{R}$ we have an orthogonal splitting
$L^{2}(\mathbb{R};\mathbb{C}^{2})=\boldsymbol{H}_{+}\oplus\boldsymbol{H}_{-},$
where $\boldsymbol{H}_{\pm}$ are the eigenspaces of the parity operator belonging to the eigenvalues $\pm 1$, respectively. Recall that the structure of $\boldsymbol{H}_{\pm}$ is the following:
$\boldsymbol{H}_{+}=\left\{\left[\begin{array}[]{c}f_{1}\\ f_{2}\end{array}\right]\in L^{2}(\mathbb{R};\mathbb{C}^{2});\,\,\text{\rm$f_{1}$ is even and $f_{2}$ is odd}\right\},$
$\boldsymbol{H}_{-}=\left\{\left[\begin{array}[]{c}f_{1}\\ f_{2}\end{array}\right]\in L^{2}(\mathbb{R};\mathbb{C}^{2});\,\,\text{\rm$f_{1}$ is odd and $f_{2}$ is even}\right\}.$
Now, Theorem 4.3 gives also a precise control on the parity of the eigenvalues of $Q_{R}$ by virtue of the control on the analytic family of eigenvectors (which are the conjugation of the eigenfunctions $u_{\pm}^{(N)}(\varepsilon)$ via the metaplectic operator $U$ in Lemma 3.6 ). We start from the case $\varepsilon=0$.
Let $v_{\pm}^{(N)}(\varepsilon):=U_{0}T^{*}u_{\pm}^{(N)}(\varepsilon)$, $|\varepsilon|<\varepsilon_{0}$, be the eigenfunctions associated with the eigenvalues $\lambda_{\pm}^{(N)}(\varepsilon)$ of $Q_{R,\varepsilon}$, respectively, which are simple if $\varepsilon\neq 0$. Also, put for short
$s_{N}:=\mathrm{sgn}(T_{+}\varphi,T_{-}\varphi)_{0}\quad\text{\rm and}\quad\sigma_{N}=(-1)^{N}s_{N}.$
Lemma 4.5.
In the notation and with the hypotheses of Theorem 4.3, the eigenvalue $\lambda_{+}^{(N)}(0)=\lambda_{-}^{(N)}(0)=N+\frac{1}{2}$ of $Q_{R,0}$ belongs both to $\mathrm{Spec}\,(Q_{R,0}|_{\boldsymbol{H}_{+}})$ and to $\mathrm{Spec}\,(Q_{R,0}|_{\boldsymbol{H}_{-}})$. In fact $v_{\pm}^{(N)}(0)\in\boldsymbol{H}_{\pm\sigma_{N}}$.
Proof.
We have that $N+\frac{1}{2}$ has multiplicity $2$ with $v_{\pm}^{(N)}(0)=U_{0}T^{*}u_{\pm}^{(N)}(0)$ generating the eigenspace of $Q_{R,0}$ associated with $N+\frac{1}{2}$ by Theorem 4.3.
If, say, $v_{+}^{(N)}(0)$ had nontrivial projections onto $\boldsymbol{H}_{+}$ and $\boldsymbol{H}_{-}$, then also $v_{+}^{(N)}(\varepsilon)$ would have nontrivial projections onto $\boldsymbol{H}_{+}$ and $\boldsymbol{H}_{+}$, for small $\varepsilon,$ whence by orthogonality the multiplicity of $\lambda_{+}^{(N)}(\varepsilon)$ would be 2 for $\varepsilon\not=0$ and small, which is impossible. Therefore, $\pm$-respectively, $v_{\pm}^{(N)}(0)$ belongs to either $\boldsymbol{H}_{+}$ or $\boldsymbol{H}_{-}$.
Next, a direct computation gives
$v_{\pm}^{(N)}(0)=\frac{1}{\|\varphi\|_{0}\sqrt{2}}\begin{bmatrix}s_{N}T_{-}\varphi\pm T_{+}\varphi\\ s_{N}T_{-}\varphi\mp T_{+}\varphi\end{bmatrix}$
and we study when, $\pm$-respectively, $v_{\pm}^{(N)}(0)$ belongs to $\boldsymbol{H}_{+}$, the case $\boldsymbol{H}_{-}$ being analogous. It is possible if and only if
$s_{N}T_{-}\varphi\pm T_{+}\varphi\quad\text{\rm is even and}\,\,s_{N}T_{-}\varphi\mp T_{+}\varphi\quad\text{\rm is odd.}$
This implies
$s_{N}\varphi(x+\alpha)=\pm\varphi(-x-\alpha),\forall x\in\mathbb{R},$
which is equivalent to
$s_{N}\varphi(x+\alpha)=\pm(-1)^{N}\varphi(x+\alpha),\forall x\in\mathbb{R},$
since an Hermite polynomial is even or odd according to whether its degree is even or odd, respectively. This proves the claim. ∎
Now we generalize the result to $\varepsilon\neq 0$.
Lemma 4.6.
In the notation and with the hypotheses of Theorem 4.3, the eigenvalue $\lambda_{\pm}^{(N)}(\varepsilon)$ belongs to $\mathrm{Spec}\,$($Q_{R,0}|_{\boldsymbol{H}_{\pm\sigma_{N}}}$)  (that is, it has parity $\pm\sigma_{N}$) for all $|\varepsilon|<\varepsilon_{0}$.
Proof.
[…]
5. The Braak conjecture
In [1], Braak conjectured that, within an interval $I_{N}:=[N,N+1),$ $N\in\mathbb{Z}_{+}$, there may be lying either $0,$ or $1$ or at most $2$ shifted eigenvalues $\lambda+g^{2}=\lambda+\alpha^{2}/2$, $\lambda\in\mathrm{Spec}(Q_{\mathrm{Rabi}})$, of the Rabi system. Moreover, if in $I_{N}$ there are two roots of the function $G_{+}$ (resp. $G_{-}$; see [1]) then the intervals $I_{N-1}$ or $I_{N+1}$ may contain zero or at most $1$ root of $G_{+}$ (analogously for $G_{-}$). Otherwise said (see [23]), Braak’s $G$-conjecture states the following: For each parity $N$ determined by the parity operator $\Pi$, the interval $I_{N}$ contains at most two shifted eigenvalues, two intervals containing no shifted eigenvalues are not adjacent, and two intervals containing two shifted eigenvalues are also not adjacent.
The latter conjecture has been proved by Rudnik in [23] ”for almost all $N$s”.
It is also convenient to restate Braak’s conjecture as in Lanuza [10]: Let $\varepsilon\not=0$ and let $B_{N}^{\pm}=\#\>\mathrm{Spec}\,(Q_{\mathrm{Rabi},\varepsilon}|_{\boldsymbol{H}_{\pm}})\cap\bigl[N+1/2,N+3/2\bigr)$. For all $N\geq 0$ and each choice of $\pm$, we have that $B_{N}^{\pm}\in\{0,1,2\}$ and $B_{N}^{\pm}+B_{N+1}^{\pm}\in\{1,2,3\}$.
In addition, as we shall see, we may state: (for $\varepsilon\not=0$ the interval $\bigl[-1/2,1/2\bigr)$ contains exactly one eigenvalue of parity $\sigma_{0}$.
Recall that for the system $Q_{R,\varepsilon}$, we have by Theorem 4.3 that for any given $j\geq 1$ there is a $k_{j}\in\mathbb{Z}_{+}$ and $c_{j}>0$ such that for the first $2(k_{j}+1)$ shifted eigenvalues of $Q_{R,\varepsilon}$ , the $\lambda^{(N)}_{\pm}(\varepsilon)$ with $N\leq k_{j},$ we have, for $|\varepsilon|<c_{j}$,
(5.1) $\lambda_{N,\pm}(\varepsilon)=N+\frac{1}{2}+\mu_{\pm}\varepsilon+o(\varepsilon),\quad N=0,1,\ldots,k_{j},$
or, recalling the definition of $\beta_{1}$ and $\beta_{2}$,
[…]
In view of Theorem 4.3, for $Q_{\mathrm{Rabi},\varepsilon}$ (and $\varepsilon\not=0$) we have the following corollary, which proves Braak’s Conjecture for finite segments of the spectrum. Because of the shift by $1/2$ in the spectrum, due to our choice of definition of $Q_{R,\varepsilon}$, the intervals in which we have to test Braak’s conjecture for the $\lambda_{\pm}^{(N)}(\varepsilon)$s are given by
$I_{N+1/2}:=\bigl[N+1/2,N+3/2\bigr),\,\,\,\,N\in\mathbb{Z}_{+}.$
Corollary 5.1 (Braak’s Conjecture).
If $4\alpha^{2}\not\in\mathsf{Z}$ and $\gamma_{1}=-\gamma_{2}$, the Braak conjecture holds for every spectral initial segment of eigenvalues: For any given $j\geq 1$ there exist $k_{j}\in\mathbb{Z}_{+}$ ($k_{j}\nearrow+\infty$) and $c_{j}>0$ sufficiently small such that the conjecture holds for the interval $I_{N+1/2}$ for all $|\varepsilon|<c_{j}$ and all $N<k_{j}-2$. More precisely, when $\varepsilon\not=0$ for each interval $\bigl[N+1/2,N+3/2\bigr)$ ($N\geq 0$) there are exactly two simple eigenvalues of $Q_{\mathrm{Rabi},\varepsilon}$ with parity $\sigma_{N}$ and $\sigma_{N+1}$, respectively. When $N=0$ (and $\varepsilon\not=0$) in $\bigl[-1/2,1/2\bigr)$ there is exactly one simple eigenvalue with parity $\sigma_{0}$.
Proof.
When $Q_{R,\varepsilon}=Q_{\mathrm{Rabi},\varepsilon}+\frac{1}{2}I_{2}$, one has $\gamma_{1}=-\gamma_{2}$ and get that $\beta_{1}=0$, $\beta_{2}=\gamma_{1}$ and, hence, $\pm\mu_{\pm}>0$ (same choice of $\pm$). Hence, for all $0\leq N<k_{j}-2$ and for $|\varepsilon|\not=0$ sufficiently small, $\lambda_{+}^{(N)}(\varepsilon),\,\lambda_{-}^{(N+1)}(\varepsilon)\in\bigl[N+1/2,N+3/2\bigr)$ and $\lambda_{+}^{(N+1)}(\varepsilon),\,\lambda_{-}^{(N+2)}(\varepsilon)\in\bigl[N+3/2,N+5/2\bigr)$ and these eigenvalues are simple by Theorem 4.3. Hence, for such $N$s, $B_{N}^{\pm}\leq 2$. Moreover, by Lemma 4.6, $\lambda_{+}^{(N)}(\varepsilon)$, $\lambda_{-}^{(N+1)}(\varepsilon)$, $\lambda_{+}^{(N+1)}(\varepsilon)$, and $\lambda_{-}^{(N+2)}(\varepsilon)$ have parity $\sigma_{N}$, $\sigma_{N+1}$, $\sigma_{N+1}$, and $\sigma_{N+2}$, respectively. Hence, as $\sigma_{N+1}$ and $\sigma_{N+1}$ have different sign, and since there are four eigenvalues in the union $\bigl[N+1/2,N+3/2\bigr)\cup\bigl[N+3/2,N+5/2\bigr)$ and they are all simple, we proved $1\leq B_{N+1}^{\pm}+B_{N+2}^{\pm}\leq 3$.
Hence, Braak’s conjecture holds. ∎
Therefore, Braak’s conjecture may fail just when we are in a degenerate case, that is when $4\alpha^{2}$ is a zero of at least a Laguerre polynomial. This case will be the topic of next section.
6. The case $4\alpha^{2}\in\mathsf{Z}$
In this section we tackle the case in which $\alpha$ is such that $4\alpha^{2}\in\mathsf{Z}$, that is, we have a family (possibly infinite) of Laguerre polynomials $L_{k}$ that vanish at $4\alpha^{2}$. So, we let
$\mathsf{K}_{\alpha}:=\{k\in\mathbb{Z}_{+};\,\,L_{k}(4\alpha^{2})=0\}.$
Hence $(T_{+}\varphi_{k},T_{-}\varphi_{k})_{0}=0$ for all $k\in\mathsf{K}_{\alpha}$ and Rellich’s approach fails in this case.
Remark 6.1.
Because of the interlacing property of the zeros of Laguerre’s polynomial we have
$k\in\mathsf{K}_{\alpha}\Longrightarrow k-1,k+1\not\in\mathsf{K}_{\alpha}.$
[…]
With $k\in\mathsf{K}_{\alpha}$ we look for
(6.2) $\lambda_{\pm,k}(\varepsilon)=\lambda_{k}+\varepsilon\lambda_{\pm,k}^{(1)}+\frac{1}{2}\varepsilon^{2}\lambda_{\pm,k}^{(2)},\quad u_{\pm,k}(\varepsilon)=\phi_{N}\otimes w_{\pm}+\varepsilon u_{\pm,k}^{(1)}+\frac{1}{2}\varepsilon^{2}u_{\pm,k}^{(2)},$
which satisfy
$(A+\varepsilon B)u_{\pm,k}(\varepsilon)-\lambda_{\pm,k}(\varepsilon)u_{\pm,k}(\varepsilon)=O(\varepsilon^{3}),$
for $|\varepsilon|$ sufficiently small (and smaller than $\varepsilon_{0}$).
Definition 6.2.
For $N\in\mathsf{K}_{\alpha}$, let $\mu_{+}=(\gamma_{1}+\gamma_{2})/2$ and let $Q^{(2)}_{N}\colon\mathbb{C}^{2}\times\mathbb{C}^{2}\longrightarrow\mathbb{C}$ be the sesquilinear form
(6.3) $Q^{(2)}_{N}(w,w^{\prime})=2\Bigl((\mu_{+}-B)R_{0}(\mu_{+}-B)\phi_{N}\otimes w,\phi_{N}\otimes w^{\prime}\Bigr)_{0},\quad w,w^{\prime}\in\mathbb{C}^{2}.$
Notice that
$Q^{(2)}_{N}(w,w^{\prime})=2\Bigl(R_{0}B\phi_{N}\otimes w,B\phi_{N}\otimes w^{\prime}\Bigr)_{0},\quad w,w^{\prime}\in\mathbb{C}^{2}.$
Theorem 6.3.
Let $N\in\mathsf{K}_{\alpha}.$ Let $Q^{(2)}_{N}$ be the bilinear form (6.3), that we suppose to possess two distinct eigenvalues $\mu_{\pm,N}^{(2)}$ (necessarily real). Let $w_{\pm}\in\mathbb{C}^{2}$ be corresponding orthonormal vectors belonging to $\mu^{(2)}_{\pm,N}$, respectively. Put in (6.2)
$\lambda_{N}=N+1/2,\,\,\lambda_{\pm,N}^{(1)}=\mu_{+},\,\,\lambda_{\pm,N}^{(2)}=\mu_{\pm,N}^{(2)},$
$u_{\pm,N}^{(1)}=R_{0}(\mu_{+}-B)\phi_{N}\otimes w_{\pm},\,\,u_{\pm,N}^{(2)}=2\mu_{\pm,N}^{(2)}R_{0}(\mu_{+}-B)R_{0}(\mu_{+}-B)\phi_{N}\otimes w_{\pm},$
[…]
Then
$(\boldsymbol{u}_{\pm,N}(\varepsilon),\lambda_{\pm,N}(\varepsilon))_{|\varepsilon|<\varepsilon_{0}}$
is a family of quasi-modes (in the sense of Lazutkin, see [11]) such that
$(A+\varepsilon B)\boldsymbol{u}_{\pm,N}(\varepsilon)-\lambda_{\pm,N}(\varepsilon)\boldsymbol{u}_{\pm,N}(\varepsilon)|\!|_{0}=O_{L^{2}}(\varepsilon^{3}),\,\,\forall\varepsilon\in I_{\varepsilon_{0}},$
where
[…]
Note that
$u_{\pm,N}(\varepsilon)|\!|_{0}=1+e_{\pm,N}(\varepsilon),\,\,|\varepsilon|\to 0+,$
where
$|e_{\pm,N}(\varepsilon)|\leq C|\varepsilon|,\,\,\,\forall\varepsilon\in I_{\varepsilon_{0}}.$
As a consequence, we have that there is $C_{1}>0$ such that
$|O_{L^{2}}(\varepsilon^{3})|\leq C_{1}^{3}|\varepsilon|^{3},\,\,\,\forall\varepsilon\in I_{\varepsilon_{0}}.$
Since
$|\lambda_{+,N}(\varepsilon)-\lambda_{-,N}(\varepsilon)|=\frac{\varepsilon^{2}}{2}|\mu_{+,N}^{(2)}-\mu_{-,N}^{(2)}|=:C_{2}^{2}\varepsilon^{2},\,\,\,\forall\varepsilon\in I_{\varepsilon_{0}},$
choosing $|\varepsilon|<\varepsilon_{0}$ such that
$2C_{1}^{3}|\varepsilon|^{3}<C_{2}^{2}\varepsilon^{2},\,\,\,\text{\rm that is}\,\,\,|\varepsilon|<C_{2}^{2}/C_{1}^{3},$
we have
$\Bigl[\lambda_{+,N}(\varepsilon)-C_{1}^{3}|\varepsilon|^{3},\lambda_{+,N}(\varepsilon)+C_{1}^{3}|\varepsilon|^{3}\Bigr]\cap\Bigl[\lambda_{-,N}(\varepsilon)-C_{1}^{3}|\varepsilon|^{3},\lambda_{-,N}(\varepsilon)+C_{1}^{3}|\varepsilon|^{3}\Bigr]=\emptyset.$
By [11], there must then be an eigenvalue of $A+\varepsilon B$ within each of the $\pm$-intervals
$\Bigl[\lambda_{\pm,N}(\varepsilon)-\varepsilon^{2},\lambda_{\pm,N}(\varepsilon)+\varepsilon^{2}\Bigr].$
But since the eigenvalues are analytic in $\varepsilon$ and the multiplicity of the eigenvalue, for $\varepsilon$ small, is at most $2$ (since the projector $\varepsilon\mapsto(2\pi i)^{-1}\int_{\Gamma}(Q_{R,\varepsilon}-z)^{-1}dz$, with $\Gamma=\{|z|=\delta\}$ contained in the resolvent set of $Q_{R,0}$, which is the projector onto the eigenspace of $N+1/2$ for $\varepsilon=0$, is holomorphic for small $\varepsilon$) we therefore have that also in this case near $N+1/2$ there are exactly $2$ distinct eigenvalues. By unitary conjugation we therefore have the following theorem.
Theorem 6.4.
Suppose $4\alpha^{2}\in\mathsf{Z}$. For each $N\in\mathsf{K}_{\alpha}$ for which the quadratic form $Q^{(2)}_{N}$ over $\mathbb{C}^{2}$ has distinct eigenvalues, we have that $Q_{R,\varepsilon}$, for $\varepsilon\not=0$ sufficiently small, has exactly $2$ eigenvalues near $N+1/2$.
When considering QRM’s spectrum near $N+1/2$ for $N\in\mathsf{K}_{\alpha}$, we have that $\mu_{+}=\mu_{-}=0,$ so that the unperturbed degenerate eigenvalue $\lambda^{(N)}_{\pm}(0)=N+1/2$ is subjected to an $\varepsilon^{2}$ perturbation. Hence, depending on whether $Q_{N}^{(2)}$ is nondegenerate and is >0>0, resp. <0<0 definite, using Remark 6.1 and the analysis of the last section, in an interval $\bigl[N+1/2,N+3/2)$ we may find, to fix ideas in case $Q_{N}^{(2)}>0$ say, the eigenvalues $\lambda_{\pm}^{(N)}(\varepsilon)$ and $\lambda_{-}^{(N+1)}(\varepsilon)$ and in this case have $0\leq B_{N}^{\pm}\leq 3$ while $0\leq B_{N-1}^{\pm}\leq 1$. However, when $Q_{N}^{(2)}$ is nondegenerate with signature $(1,1)$, one may again have a configuration of the spectrum consistent with Braak’s conjecture.
The point is to control the parity of the eigenfunctions near the eigenvalue $N+1/2$ of the unperturbed operator.
Lemma 6.5.
For all $N\geq 0$ and for any choice of $\pm$, $\Pi_{\boldsymbol{H}_{\pm}}\mathrm{Ker}(Q_{R,0}-(N+1/2))\neq\{0\}$ where $\Pi_{\boldsymbol{H}_{\pm}}$ is the projection onto $\boldsymbol{H}_{\pm}$.
Proof.
In the first place, we know that
(6.4) $\mathrm{Ker}(Q_{R,0}-(N+1/2))=\mathrm{Span}\,\{U_{0}T^{*}(\phi_{N}\otimes e_{\pm})\}=\mathrm{Span}\,\Big\{(T_{\mp}\phi_{N})\otimes\begin{bmatrix}1\\ \pm 1\end{bmatrix}\Big\}$,
where $e_{+}=\left[\begin{array}[]{c}1\\ 0\end{array}\right]$ and $e_{-}=\left[\begin{array}[]{c}0\\ 1\end{array}\right]$, respectively. Next, since $\boldsymbol{H}_{+}\oplus\boldsymbol{H}_{-}=L^{2}$, it is sufficient to show that $\mathrm{Ker}(Q_{R,0}-(N+1/2))\not\subset\boldsymbol{H}_{\pm}$ to prove the claim. In fact, by (6.4) having $\mathrm{Ker}(Q_{R,0}-(N+1/2))\subset\boldsymbol{H}_{+}$ is equivalent to $T_{\mp}\phi_{N}$ is even and $\pm T_{\mp}\phi_{N}$ is odd. However, this is impossible since the only function being both even and odd is 0. The same argument applies in case of inclusion in $\boldsymbol{H}_{-}$. This concludes the proof. ∎
The foregoing lemma allows us to obtain a result on the parity of the two distinct eigenvalues in Theorem 6.4 when $\varepsilon\neq 0$, as the next result shows.
Lemma 6.6.
In the notation of Theorem 6.4, the two distinct eigenvalues near $N+1/2$ have different parity for $\varepsilon\neq 0$.
Proof.
Let $\varepsilon\mapsto v_{\pm}^{(N)}(\varepsilon)$ be the analytic family of eigenfunctions of $Q_{R,\varepsilon}$ associated with the eigenvalues $\lambda_{\pm}^{(N)}(\varepsilon)$ of $Q_{R,\varepsilon}$.
As a first step, we show that either $v_{+}^{(N)}(\varepsilon)\in\boldsymbol{H}_{+}$ or $v_{+}^{(N)}(\varepsilon)\in\boldsymbol{H}_{-}$ (we can proceed similarly for $v_{-}^{(N)}(\varepsilon)$) for all $\varepsilon\neq 0$. In fact, $\Pi_{\boldsymbol{H}_{\pm}}v_{+}^{(N)}(\varepsilon)$ also satisfy the eigenvalue equation with respect to $\lambda_{+}^{(N)}(\varepsilon),$ since $Q_{R,\varepsilon}=Q_{\mathrm{Rabi},\varepsilon}$ (recall that $\gamma_{1}=-\gamma_{2}$) commutes with the parity operator. If both projections were nonzero, they would be eigenfunctions as well, and because of their orthogonality they would be linearly independent, yielding that $\lambda_{+}^{(N)}(\varepsilon)$ could not be simple for $\varepsilon\not=0$.
Next, as a second step, we show that if $v_{+}^{(N)}(\varepsilon)\in\boldsymbol{H}_{\sigma}$ ($\sigma=1$ or 1-1) for some $\varepsilon$, then $v_{+}^{(N)}(\varepsilon)\in\boldsymbol{H}_{\sigma}$ for all $\varepsilon$ (we can proceed similarly for $v_{-}^{(N)}$). Assume, to fix ideas, $v_{+}^{(N)}(\varepsilon^{\prime})\in\boldsymbol{H}_{+}$. If there were $\varepsilon^{\prime}\not=0$ and $\delta>0$ such that $v_{+}^{(N)}(\varepsilon^{\prime})\in\boldsymbol{H}_{+}$ for $\varepsilon\in(\varepsilon^{\prime}-\delta,\varepsilon^{\prime})$ and $v_{+}^{(N)}(\varepsilon)\in\boldsymbol{H}_{-}$ for $\varepsilon\in(\varepsilon^{\prime},\varepsilon^{\prime}+\delta)$, then by continuity of the family $\varepsilon\mapsto v_{+}^{(N)}(\varepsilon)$ we would have $0\neq v_{+}^{(N)}(\varepsilon^{\prime})\in\boldsymbol{H}_{+}\cap\boldsymbol{H}_{-}=\{0\}$ (by the orthogonality of $\boldsymbol{H}_{+}$ and $\boldsymbol{H}_{-}$).
As a final step, we show that the two eigenvalues $v_{\pm}^{(N)}(\varepsilon)$ cannot be in the same space $\boldsymbol{H}_{+}$ or $\boldsymbol{H}_{-}$ for any fixed $\varepsilon$. In fact, if by contradiction we had that both eigenfunctions $v_{\pm}^{(N)}(\varepsilon)\in\boldsymbol{H}_{+}$ for some $\varepsilon$ (we can proceed similarly for $\boldsymbol{H}_{-}$), then, by the second step, we would have $v_{\pm}^{(N)}(0)\in\boldsymbol{H}_{+}$ and, hence, $\Pi_{\boldsymbol{H}_{-}}\mathrm{Ker}(Q_{R,0}-(N+1/2))=\Pi_{\boldsymbol{H}_{-}}\mathrm{Span}\,\{v_{\pm}^{(N)}(0)\}=\{0\}$ against Lemma 6.5. This concludes the proof. ∎
We may therefore state the following corollary of our approach.
Corollary 6.7 (Braak’s conjecture, degenerate case).
Consider our system $Q_{R,\varepsilon}$ in the case of $Q_{\mathrm{Rabi},\varepsilon}$, i.e. when $\mu_{\pm}=0$. Suppose the bilinear form $Q_{N}^{(2)}$ is nondegenerate with signature $(1,1).$ Then Braak’s conjecture keeps holding true.
It will be interesting also to consider the other cases depending on the positivity of $Q_{N}^{(2)}$, and also cases in which the eigenvalues are either constant (i.e. $N+1/2$ for some $N\in\mathsf{K}_{\alpha}$) or have derivatives with respect to $\varepsilon$ at 0 all zero up to some $k_{0}\in\mathbb{N}$ (hence $d^{k_{0}+1}\lambda_{\pm}^{(N)}/d\varepsilon^{k_{0}+1}(0)\not=0$). In that case the behavior will be decided (using, for instance, a quasi-mode construction) by an analogous bilinear form similar to $Q_{N}^{(2)}$ in the case $k_{0}=1$.
6.1. An explicit expression for the form $Q_{N}^{(2)}$.
In this subsection we wish to show how it is possible to explicitly write down the quadratic form $Q_{N}^{(2)}.$ It shows that a combination of Rellich’s formulae and a quasimode construction gives the possibility to study highly degenerate eigenvalues. Since as functions of $\varepsilon$ the eigenvalues are analytic, this gives the opportunity to understand the spectrum (of course, in the perturbative regime) of the QRM.
Recall that, with $\lambda_{k}=k+1/2$ for $k\in\mathbb{Z}_{+},$
$R_{0}(\mu_{+}-B)\phi_{N}\otimes w=-\sum_{j=\pm}\sum_{k\not=N}\frac{1}{\lambda_{k}-\lambda_{N}}\bigl(B\phi_{N}\otimes w,\phi_{k}\otimes w_{j}\bigr)_{0}\phi_{k}\otimes w_{j},$
for all $w\in\mathbb{C}^{2}$, where $\{w_{+},w_{-}\}$ is the orthonormal basis of $\mathbb{C}^{2}$ chosen above. It follows that
$Q_{N}^{(2)}(w,w^{\prime})=\sum_{j=\pm}\sum_{k\not=N}(\lambda_{k}-\lambda_{N})^{-1}\bigl(B\phi_{N}\otimes w,\phi_{k}\otimes w_{j}\bigr)_{0}\overline{\bigl(B\phi_{N}\otimes w^{\prime},\phi_{k}\otimes w_{j}\bigr)_{0}}.$
Let
$M_{N,k}=\left[\begin{array}[]{cc}0&(T_{+}\phi_{N},T_{-}\phi_{k})_{0}\\ (T_{-}\phi_{N},T_{+}\phi_{k})_{0}&0\end{array}\right].$
We may then write
$(B\phi_{N}\otimes w,\phi_{k}\otimes w_{\pm})_{0}=\langle M_{N,k}w,w_{\pm}\rangle_{\mathbb{C}^{2}},\quad w\in\mathbb{C}^{2}.$
Therefore, writing for $\zeta,$ $w,$ $w^{\prime}\in\mathbb{C}^{2}$ the action of the tensor-product
$\bar{\zeta}\otimes\zeta(w)=\langle w,\zeta\rangle_{\mathbb{C}^{2}}\zeta,$
so that
$\langle(\bar{\zeta}\otimes\zeta)w,w^{\prime}\rangle_{\mathbb{C}^{2}}=\langle w,\zeta\rangle_{\mathbb{C}^{2}}\,\overline{\langle w^{\prime},\zeta\rangle_{\mathbb{C}^{2}}},$
we obtain the following formula
$Q_{N}^{(2)}(w,w^{\prime})=2\Bigl\langle\Bigl(\sum_{j=\pm}\sum_{k\not=N}\frac{\overline{M^{*}_{N,k}w_{j}}\otimes M_{N,k}^{*}w_{j}}{\lambda_{k}-\lambda_{N}}\Bigr)w,w^{\prime}\Bigr\rangle_{\mathbb{C}^{2}}.$
7. The Weyl law for some generalizations of the Quantum Rabi Model
In this section we deal with the asymptotic spectral properties of some generalizations of the Quantum Rabi Model. We extend to this setting the $2$-term asymptotics of the Weyl spectral counting function obtained in Malagutti and Parmeggiani [13] for semiregular SMGES systems (see Definition 3.5 above) modelled after the Jaynes-Cummings Model. In that paper the Jaynes-Cummings Model is written as a semiregular Non-Commutative Harmonic Oscillator, in the SMGES class as introduced in [13]: studying the high-energy spectral properties of this class of pseudodifferential systems, a Weyl law and a refined Weyl law (Theorems 7.8 and 7.9 in [13]) were obtained. (For the same class, also the meromorphic continuation of the spectral zeta-function [12], and a spectral quasi-clustering property [14] have been obtained.) We next show how to obtain a Weyl law for a class of systems that includes and generalizes the QRM.
As already remarked, QRM is not an SMGES, in that, although the system is principally scalar, the semiprincipal symbol cannot be smoothly diagonalized in the class $S^{m}(\mathbb{R}^{n};\mathsf{M}_{N})$. However, we may construct a perturbation that will be used to find spectral asymptotics of the class of systems we are interested in.
7.1. Quantum Rabi Model generalizations
In this section we introduce a class of semiregular NCHOs which generalizes the Quantum Rabi Model. For this class we state and prove a Weyl law result in Section 7.2 below. Let $E_{jk}$ be the $N\times N$ matrix which is $1$ in the $jk$ entry and 0 otherwise. In what follows $p_{2}^{\mathrm{w}}(x,D)=(|D|^{2}+|x|^{2})/2$ will be the $n$-dimensional quantum harmonic oscillator. One has the following list of models that generalize the QRM, in the different instances in analogy with the Jaynes-Cummings systems [8].
7.1.1. The Quantum Rabi Model for an $N$-level atom and n=N1n=N-1 cavity-modes in the $\Xi$-configuration
In this case, for $\alpha_{1},\ldots\alpha_{N-1}\in\mathbb{R}\setminus\{0\}$, $\gamma_{1},\ldots\gamma_{N-1}\in\mathbb{R}$ with $\gamma_{1}\leq\gamma_{2}\leq\ldots\leq\gamma_{N-1}$, we consider the $N\times N$ system in $\mathbb{R}^{n}$, n=N1n=N-1, given by
$a^{\mathrm{w}}(x,D)=p_{2}^{\mathrm{w}}(x,D)I_{N}+\sum_{k=1}^{N-1}\alpha_{k}x_{k}\Bigl(E_{k,k+1}+E_{k+1,k}\Bigr)+\sum_{k=1}^{N-1}\gamma_{k}E_{k+1,k+1}.$
In this case, the levels of the atom are given by 0 and the $\gamma_{k}.$
7.1.2. The Quantum Rabi Model for an $N$-level atom and n=N1n=N-1 cavity-modes in the $\bigwedge$-configuration
In this case, for $\alpha_{1},\ldots\alpha_{N-1}\in\mathbb{R}\setminus\{0\}$, $\gamma_{1},\ldots\gamma_{N-1}\in\mathbb{R}$ with $\gamma_{1}\leq\gamma_{2}\leq\ldots\leq\gamma_{N-1}$, we consider the $N\times N$ system in $\mathbb{R}^{n}$, n=N1n=N-1, given by
$a^{\mathrm{w}}(x,D)=p_{2}^{\mathrm{w}}(x,D)I_{N}+\sum_{k=1}^{N-1}\alpha_{k}x_{k}\Bigl(E_{k,N}+E_{N,k}\Bigr)+\sum_{k=1}^{N-1}\gamma_{k}E_{k+1,k+1}.$
In this case, the levels of the atom are given by 0 and the $\gamma_{k}.$
7.1.3. The Quantum Rabi Model for an $N$-level atom and n=N1n=N-1 cavity-modes in the so-called $\bigvee$-configuration
In this case, for $\alpha_{1},\ldots\alpha_{N-1}\in\mathbb{R}\setminus\{0\}$, $\gamma_{1},\ldots\gamma_{N-1}\in\mathbb{R}$ with $\gamma_{1}\leq\gamma_{2}\leq\ldots\leq\gamma_{N-1}$, we consider the $N\times N$ system in $\mathbb{R}^{n}$, n=N1n=N-1, given by
$a^{\mathrm{w}}(x,D)=p_{2}^{\mathrm{w}}(x,D)I_{N}+\sum_{k=1}^{N-1}\alpha_{k}x_{k}\Bigl(E_{1,k+1}+E_{k+1,1}\Bigr)+\sum_{k=1}^{N-1}\gamma_{k}E_{k+1,k+1}.$
In this case, the levels of the atom are given by 0 and the $\gamma_{k}.$
7.2. A Weyl law for the generalized Quantum Rabi Models
We may state the following Weyl law.
Theorem 7.1.
Let $a=a^{*}\in S^{2}_{\mathrm{sreg}}(\mathbb{R}^{n};\mathsf{M}_{N})$ with
$a\sim\sum_{j\geq 0}a_{2-j},$
(n=N1n=N-1) where $a_{k}=a_{k}^{*}$ is positively homogeneous of degree $k$. Suppose that $a_{2}=p_{2}I_{N}$, where now $p_{2}(X)=|X|^{2}/2,$ $X\in\mathbb{R}^{2n}$, is the $n$-d scalar harmonic oscillator and that on $\dot{\mathbb{R}}^{2n}$ there is a smooth matrix-valued function $b_{1}$, positively homogeneous of degree $1$, such that $a_{\epsilon}:=a+\varepsilon b_{1}$ is an SMGES for all $\varepsilon\in(0,1)$. Suppose that $A=a^{\mathrm{w}}(x,D)$, realized as an unbounded operator in $L^{2}$ with maximal domain $B^{2}$, is positive.
Then, with $\mathbb{R}\ni\lambda\longmapsto\mathsf{N}_{A}(\lambda)$ denoting the spectral Weyl counting function associated with $A$,
$\mathsf{N}_{A}(\lambda)=(2\pi)^{-n}\left(N\lambda^{n}\int_{p_{2}\leq 1}dX-\lambda^{n-1/2}\int_{p_{2}=1}\mathsf{Tr}(a_{1})\frac{ds}{|\nabla p_{2}|}\right)+o(\lambda^{n-1/2}),$
as $\lambda\rightarrow+\infty$. (Here $ds$ is the Riemannian measure induced on $\{p_{2}=1\}$ by the $2n$-dimensional Euclidean one.)
Proof.
To prove the theorem we use a perturbation argument. We consider $A=a^{\mathrm{w}}(x,D)$ realized as an unbounded operator in $L^{2}(\mathbb{R}^{n};\mathbb{C}^{N})$ with maximal domain $D(A)=B^{2}(\mathbb{R}^{n};\mathbb{C}^{N})$, and write $B_{1}=b_{1}^{\mathrm{w}}(x,D)$ realized also as an unbounded operator in $L^{2}$ with domain $B^{1}(\mathbb{R}^{n};\mathbb{C}^{N})$. We then let $A_{\varepsilon}=A+\varepsilon B_{1}$ realized on the domain $D(A_{\varepsilon})=D(A)$ (independent of $\varepsilon$). The point will be to obtain an operator inequality between $A_{\varepsilon}$ and $A$ which, by the minimax principle, leads to a spectral inequality between $A_{\varepsilon}$ and $A$. The perturbation $A_{\varepsilon}$ belongs to the class SMGES for $\varepsilon>0$. Next, we use Theorem 7.8 (Weyl law) in [13] to have a Weyl law for $A_{\varepsilon}$. Finally, we can obtain a Weyl law for $A$ by the spectral inequality just proven. Actually, the Weyl law for the second order operator $A$ is not refined since the perturbation has order $1$, whence only the first term after the leading term of the asymptotics can be determined precisely.
[…]
$|(A^{-1/4}(A_{\varepsilon}-A)A^{-1/4}\phi,\phi)_{L^{2}}|=$ $\varepsilon|(A^{-1/4}B_{1}A^{-1/4}\phi,\phi)_{L^{2}}|$
[…]
(7.2) $(A\psi,\psi)_{L^{2}}-\varepsilon C(A^{1/2}\psi,\psi)_{L^{2}}\leq(A_{\varepsilon}\psi,\psi)_{L^{2}}\leq(A\psi,\psi)_{L^{2}}+\varepsilon C(A^{1/2}\psi,\psi)_{L^{2}}.$
Let $0<\lambda_{1}\leq\lambda_{2}\leq\ldots$ (respectively, $\lambda_{1,\varepsilon}\leq\lambda_{2,\varepsilon}\leq\ldots$) be the eigenvalues of $A$ (respectively $A_{\varepsilon}$), repeated according to multiplicities. By the minimax principle and (7.2) we have
$\lambda_{j}-\varepsilon C\sqrt{\lambda_{j}}\leq\lambda_{j,\varepsilon}\leq\lambda_{j}+\varepsilon C\sqrt{\lambda_{j}},\quad\forall j\geq 1,$
which leads to an estimate for the counting function $\mathsf{N}_{A}$ of $A$ and $\mathsf{N}_{A_{\varepsilon}}$ of $A_{\varepsilon}$. In fact, for $\varepsilon$ sufficiently small
$\mathsf{N}_{A_{\varepsilon}}(\lambda):=$ $\#\{j;\,\,\lambda_{j,\varepsilon}\leq\lambda\}$
$\geq$ $\#\{j;\,\,\lambda_{j}+\varepsilon C\sqrt{\lambda_{j}}\leq\lambda\}$
$=$ $\#\{j;\,\,\nu_{\varepsilon,+}^{-1}(\lambda_{j})\leq\lambda\}=\mathsf{N}_{A}(\nu_{\varepsilon,+}(\lambda)),$
and
$\mathsf{N}_{A_{\varepsilon}}(\lambda):=$ $\#\{j;\,\,\lambda_{j,\varepsilon}\leq\lambda\}$
$\leq$ $\#\{j;\,\,\lambda_{j}-\varepsilon C\sqrt{\lambda_{j}}\leq\lambda\}$
$=$ $\#\{j;\,\,\nu_{\varepsilon,-}^{-1}(\lambda_{j})\leq\lambda\}=\mathsf{N}_{A}(\nu_{\varepsilon,-}(\lambda)),$
where, with $c_{0}>0$ a lower bound of $\mathrm{Spec}(A)$,
$\nu_{\varepsilon,\pm}:(c_{0},+\infty)\longrightarrow(\nu_{\varepsilon,\pm}(c_{0}),+\infty),\quad\lambda\mapsto\frac{\varepsilon^{2}}{2}C^{2}+\lambda\mp\varepsilon C\sqrt{\frac{\varepsilon^{2}}{4}C^{2}+\lambda},$
is a smooth function which is strictly increasing and invertible for $\varepsilon<\varepsilon_{0}:=2\sqrt{c_{0}}/C,$ with inverse given by
$\nu_{\varepsilon,\pm}^{-1}:\lambda\mapsto\lambda\pm\varepsilon C\sqrt{\lambda}.$
Therefore, for $\varepsilon<\varepsilon_{0}$
$\mathsf{N}_{A_{\varepsilon}}\bigl(\nu_{\varepsilon,-}^{-1}(\lambda)\bigr)\leq\mathsf{N}_{A}(\lambda)\leq\mathsf{N}_{A_{\varepsilon}}\bigl(\nu_{\varepsilon,+}^{-1}(\lambda)\bigr),$
which is equivalent to
$\frac{1}{\lambda^{n-1/2}}\left(\mathsf{N}_{A_{\varepsilon}}\bigl(\nu_{\varepsilon,-}^{-1}(\lambda)\bigr)-(2\pi)^{-n}\lambda^{n}N\int_{p_{2}\leq 1}dX\right)$
$\leq$ $\frac{1}{\lambda^{n-1/2}}\left(\mathsf{N}_{A}(\lambda)-(2\pi)^{-n}\lambda^{n}N\int_{p_{2}\leq 1}dX\right)$
(7.3) $\leq$ $\frac{1}{\lambda^{n-1/2}}\left(\mathsf{N}_{A_{\varepsilon}}\bigl(\nu_{\varepsilon,+}^{-1}(\lambda)\bigr)-(2\pi)^{-n}\lambda^{n}N\int_{p_{2}\leq 1}dX\right).$
Now, we study the behavior of $\mathsf{N}_{A_{\varepsilon}}\circ\nu_{\varepsilon,+}^{-1}$ when $\lambda\rightarrow+\infty$. By Theorem 7.8 in [13] and since the semiprincipal symbol of $A_{\varepsilon}$ is $a_{1}+\varepsilon b_{1}$,
$\mathsf{N}_{A_{\varepsilon}}\bigl(\nu_{\varepsilon,+}^{-1}(\lambda)\bigr)=$ $(2\pi)^{-n}\biggl(N\,\nu_{\varepsilon,+}^{-1}(\lambda)^{n}\int_{p_{2}\leq 1}dX$
$-\nu_{\varepsilon,+}^{-1}(\lambda)^{n-1/2}\int_{p_{2}=1}\mathsf{Tr}(a_{1}+\varepsilon b_{1})\frac{ds}{|\nabla p_{2}|}\biggr)+O_{\varepsilon}(\lambda^{n-1})$
$=$ $(2\pi)^{-n}\left(\lambda^{n}N\left(\int_{p_{2}\leq 1}dX\right)\right$.
$-\lambda^{n-1/2}\biggl(\int_{p_{2}=1}{\rm\mathsf{Tr}}(a_{1}+\varepsilon b_{1})\frac{ds}{|\nabla p_{2}|}-\varepsilon nCN\int_{p_{2}\leq 1}dX\biggr)\biggr)+O_{\varepsilon}(\lambda^{n-1}),$
as $\lambda\rightarrow+\infty$, where the constant in $O_{\varepsilon}(\lambda)$ may depend on $\varepsilon$. In a similar way, for $\mathsf{N}_{A_{\varepsilon}}\circ\nu_{\varepsilon,-}^{-1}$ we have
$\mathsf{N}_{A_{\varepsilon}}(\nu_{\varepsilon,-}^{-1}(\lambda))=$ $(2\pi)^{-n}\left(\lambda^{n}N\left(\int_{p_{2}\leq 1}dX\right)\right$.
$-\lambda^{n-1/2}\biggl(\int_{p_{2}=1}{\rm\mathsf{Tr}}(a_{1}+\varepsilon b_{1})\frac{ds}{|\nabla p_{2}|}+\varepsilon nCN\int_{p_{2}\leq 1}dX\biggr)\biggr)+O_{\varepsilon}(\lambda^{n-1}),$
as $\lambda\rightarrow+\infty$.
Hence, since we are looking for the behavior of $\mathsf{N}_{A}$ as $\lambda\rightarrow+\infty$, and since one has $O_{\varepsilon}(\lambda^{n-1})/\lambda^{n-1/2}\xrightarrow[\lambda\to+\infty]{}0$ for all $\varepsilon$, by (7.3) we get
$-(2\pi)^{-n}\left(\int_{p_{2}=1}{\rm\mathsf{Tr}}(a_{1}+\varepsilon b_{1})\frac{ds}{|\nabla p_{2}|}+\varepsilon nCN\int_{p_{2}\leq 1}dX\right)$
$\leq$ ${\displaystyle\liminf_{\lambda\rightarrow+\infty}}\left(\frac{1}{\lambda^{n-1/2}}\left(\mathsf{N}_{A}(\lambda)-(2\pi)^{-n}\lambda^{n}N\int_{p_{2}\leq 1}dX\right)\right)$
$\leq$ ${\displaystyle\limsup_{\lambda\rightarrow+\infty}}\left(\frac{1}{\lambda^{n-1/2}}\left(\mathsf{N}_{A}(\lambda)-(2\pi)^{-n}\lambda^{n}N\int_{p_{2}\leq 1}dX\right)\right)$
$\leq$ $-(2\pi)^{-n}\left(\int_{p_{2}=1}{\rm\mathsf{Tr}}(a_{1}+\varepsilon b_{1})\frac{ds}{|\nabla p_{2}|}-\varepsilon nCN\int_{p_{2}\leq 1}dX\right),$
for all $\varepsilon<\varepsilon_{0}$ which implies, by taking the limit as $\varepsilon\rightarrow 0+$,
$-(2\pi)^{-n}\int_{p_{2}=1}{\rm\mathsf{Tr}}(a_{1})\frac{ds}{|\nabla p_{2}|}$ $\leq{\displaystyle\liminf_{\lambda\rightarrow+\infty}}\left(\frac{1}{\lambda^{n-1/2}}\left(\mathsf{N}_{A}(\lambda)-(2\pi)^{-n}\lambda^{n}N\int_{p_{2}\leq 1}dX\right)\right)$
$\leq{\displaystyle\limsup_{\lambda\rightarrow+\infty}}\left(\frac{1}{\lambda^{n-1/2}}\left(\mathsf{N}_{A}(\lambda)-(2\pi)^{-n}\lambda^{n}N\int_{p_{2}\leq 1}dX\right)\right)$
$\leq-(2\pi)^{-n}\int_{p_{2}=1}{\rm\mathsf{Tr}}(a_{1})\frac{ds}{|\nabla p_{2}|},$
and completes the proof. ∎
Theorem 7.1 may be used with n=N1n=N-1 to obtain a non-refined Weyl law for the models of Subsection 7.1. Namely, in the notation of Theorem 7.1:
  • For the models of Subsection 7.1.1 we have that
    $a=p_{2}I_{N}+\sum_{k=1}^{N-1}\alpha_{k}x_{k}\Bigl(E_{k,k+1}+E_{k+1,k}\Bigr)+\sum_{k=1}^{N-1}\gamma_{k}E_{k+1,k+1},$
    $b_{1}=\sum_{k=1}^{N-1}\alpha_{k}\Bigl((i\xi_{k})^{*}E_{k,k+1}+i\xi_{k}E_{k+1,k}\Bigr).$
    In fact,
    $a+\varepsilon b_{1}=$ $p_{2}I_{N}+\sum_{k=1}^{N-1}\sqrt{2}\alpha_{k}\Bigl(\psi_{k,\varepsilon}^{*}E_{k,k+1}+\psi_{k,\varepsilon}E_{k+1,k}\Bigr)$
    $+\sum_{k=1}^{N-1}\gamma_{k}E_{k+1,k+1},$
    where $\psi_{k,\varepsilon}:=(x_{k}+i\varepsilon\xi_{k})/\sqrt{2},$ $1\leq k\leq n,$ is an SMGES for all $\varepsilon>0$;
  • For the models of Subsection 7.1.2, Theorem 7.1 can be used to obtain a non-refined Weyl law with
    $a=p_{2}I_{N}+\sum_{k=1}^{N-1}\alpha_{k}x_{k}\Bigl(E_{k,N}+E_{N,k}\Bigr)+\sum_{k=1}^{N-1}\gamma_{k}E_{k+1,k+1},$
    $b_{1}=\sum_{k=1}^{N-1}\alpha_{k}\Bigl((i\xi_{k})^{*}E_{k,N}^{*}+i\xi_{k}E_{N,k}\Bigr),$
    while for the models of Subsection 7.1.2, Theorem 7.1 can be used with
    $a=p_{2}I_{N}+\sum_{k=1}^{N-1}\alpha_{k}x_{k}\Bigl(E_{1,k+1}+E_{k+1,1}\Bigr)+\sum_{k=1}^{N-1}\gamma_{k}E_{k+1,k+1},$
    $b_{1}=\sum_{k=1}^{N-1}\alpha_{k}\Bigl((i\xi_{k})^{*}E_{k,k+1}^{*}+i\xi_{k}E_{k+1,k}\Bigr).$
Data Availability Statement. All data is provided in full in this paper.
Declarations
Conflict of interest.
There is no conflict of interest.